OpenStudy (anonymous):

A store is having a sale on jelly beans and trail mix. For 2 pounds of jelly beans and 5 pounds of trail mix, the total cost is $10 . For 8 pounds of jelly beans and 3 pounds of trail mix, the total cost is $23 . Find the cost for each pound of jelly beans and each pound of trail mix.

OpenStudy (anonymous):

So could U show me your steps step by step just as I mentioned before ? Hint : Let abbreviations to be (JM) for Jelly bean and (TM) from trail matrix :D Step by step and don't worry it's very simple as ones before :)

OpenStudy (anonymous):

@Omar_Elboredy could you go step by step for me

OpenStudy (mathstudent55):

Let the two unknown unit prices be variables. Let the unit price of jelly beans be b, and let the unit price of trail mix be m.

OpenStudy (mathstudent55):

For 2 pounds of jelly beans and 5 pounds of trail mix, the total cost is $10. If the unit price of jelly beans is b, and you buy 2 pounds, what is the cost of the jelly beans?

OpenStudy (anonymous):

$5 @mathstudent55

OpenStudy (mathstudent55):

@Deldrickg Are you there?

OpenStudy (anonymous):

yees just responded

OpenStudy (anonymous):

So from the problem statement and abbreviations I told to U : 2JB + 5 TM = 10 & 8JB + 3TM = 23 nw multiply (eq.1) by 4 U will get : 8JB + 20TM = 40 and then substract (eq.2) from the result eq ... U will get then : 17TM = 17 , then TM = 17/17 = 1 $ And finally U substitute back by TM = 1 in any of the equations (eq.1) or (eq.2) ,U will get : JB = 2.5$

OpenStudy (mathstudent55):

2 pounds of jelly beans costs 2b 5 pounds of trail mix costs 5m The total is $10, so you have your first equation: 2b + 5m = 10

OpenStudy (mathstudent55):

You do the same for the second equation, and you get: For 8 pounds of jelly beans and 3 pounds of trail mix, the total cost is $23. 8b + 3m = 23 This is your system of equations:: 2b + 5m = 10 8b + 3m = 23

OpenStudy (mathstudent55):

Multiply the first equation by -4 on both sides and add the equations: -8b - 20m = -40 8b + 3m = 23 ---------------- -17m = -17 Divide both sides by -17 to get m = 1 Now substitute 1 for m in the first equation: 2b + 5m = 10 2b + 5(1) = 10 2b + 5 = 10 2b = 5 b = 2.5 Answer: The jelly beans cost $2.50 per pound, and the trail mix costs $1 per pound.